I am having solving systems through elimination trouble!
Here are the equations:
0.3x-0.2y=4
0.4x+0.3y=5/19
I can solve for "y" but when I get here: (I am lost)
3x-2(-170/19)=40 the answer:140/19 (these are fractions a/b)
Please help me

Answer by Alan3354(69443)

0.3x-0.2y=4
0.4x+0.3y=5/19
----------------
3x - 2y = 40 (x 10)
76x + 57y = 50 (x 190)
---------------------
171x - 114x = 2280 (x 57)
152x + 114y = 100 (x 2)
------------------------ Add
323x = 2380
x = 2380/323
x = 140/19
-------------
3*(140/19) - 2y = 40
-2y = 340/19
y = -170/19